OpenStudy (anonymous):

evaluate the integral from 0 to pi/2 [(2-sinx)^2dx]

OpenStudy (raden):

Fisrt, u have to expand this : (2-sinx)^2 = 4 - 4sin(x) + sin^2 (x)

OpenStudy (raden):

use the identity to converting sin^2 (x) = (1 - cos(2x))/2 = 1/2 - 1/2*cos(2x)

OpenStudy (raden):

so, (2-sinx)^2 = 4 - 4sin(x) + 1/2 - 1/2*cos(2x) = 9/2 - 4sin(x)- 1/2*cos(2x) now, u can integrate each term one by one just use the basic formula in integration

OpenStudy (anonymous):

ohhhh got it. thanks :)

OpenStudy (raden):

you're welcome :)